OpenStudy (elenathehomeschooler):

The number of people in a town of 10,000 who have heard a rumor started by a small group of people is given by the following function: N(t) = 10,000/(5 + 1245e^-.97t) 15. How many people have heard the rumor after 1 day? After 5 days? 16. How long before 1000 of the people in the town have heard the rumor?

OpenStudy (elenathehomeschooler):

@jim_thompson5910 can you help?

jimthompson5910 (jim_thompson5910):

what did you get when you plugged in t = 1 ?

OpenStudy (elenathehomeschooler):

i got 2, i dont think i got it right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's too small. How did you type it into your calc?

OpenStudy (elenathehomeschooler):

N(1)=10000/5+1245e^-.97(1)

jimthompson5910 (jim_thompson5910):

N(t) = 10000/(5+1245*e^(-0.97*t)) N(1) = 10000/(5+1245*e^(-0.97*1)) ... replace t with 1 so you should type in `10000/(5+1245*e^(-0.97*1))` into your calc

OpenStudy (elenathehomeschooler):

20.9661898567944051

jimthompson5910 (jim_thompson5910):

yep, round that to the nearest whole number

jimthompson5910 (jim_thompson5910):

now repeat what you just did, but do it for t = 5 now

OpenStudy (elenathehomeschooler):

21 people in 1 day and 678 people in 5 days

OpenStudy (elenathehomeschooler):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

for #16, you'll need to solve the equation \[\Large 1,000 = \frac{10,000}{5+1245e^{-0.97t}}\] for t

OpenStudy (elenathehomeschooler):

i got 5.6876 rounded it would be 6 days @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (elenathehomeschooler):

Great! thanks!

jimthompson5910 (jim_thompson5910):

you're welcome